# Cog-TiPRO: Iterative Prompt Refinement with LLMs to Detect Cognitive Decline via Longitudinal Voice Assistant Commands

### v2

@@ Abstract
Early detection of cognitive decline is crucial for enabling interventions that can slow neurodegenerative disease progression. Traditional diagnostic approaches rely on labor-intensive clinical assessments, which are impractical for frequent monitoring. Our pilot study investigates voice assistant systems (VAS) as non-invasive tools for detecting cognitive decline through longitudinal analysis of speech patterns in voice commands. Over an 18-month period, we collected voice commands from 35 older adults, with 15 participants providing daily at-home VAS interactions. To address the challenges of analyzing these short, unstructured and noisy commands, we propose Cog-TiPRO, a framework that combines (1) LLM-driven iterative prompt refinement for linguistic feature extraction, (2) HuBERT-based acoustic feature extraction, and (3) transformer-based temporal modeling. Using iTransformer, our approach achieves 73.80% accuracy and 72.67% F1-score in detecting MCI, outperforming its baseline by 27.13%. Through our LLM approach, we identify linguistic features that uniquely characterize everyday command usage patterns in individuals experiencing cognitive decline.
@@
Mild Cognitive Impairment (MCI), a precursor to dementia, is characterized by subtle declines in cognitive functions such as attention, memory, and language. Early detection of these changes enables timely intervention to mitigate progression to dementia. However, detection and continuous monitoring of MCI remain challenging due to the absence of gold standard biomarkers. Additionally, determining whether cognitive changes represent pathological decline requires complex clinical judgment. Currently, cognitive impairment often remains undetected until significant progression to dementia has occurred, leaving patients and families unprepared for managing its cognitive and functional impacts [1].
Traditional diagnostics for MCI rely on clinical assessments, including neuroimaging, laboratory tests, and standardized neuropsychological examinations such as the Montreal Cognitive Assessment (MoCA) [2]. While effective for diagnosis, these methods have limitations: They require administration by trained healthcare professionals in clinical settings, making them labor-intensive and expensive for frequent monitoring. Since speech impairments such as increased pauses and language disfluency are among the earliest symptoms of cognitive decline, speech processing represents a promising approach for non-invasive and cost-effective MCI detection that could potentially be implemented in everyday environments [3].
Prior research in cognitive monitoring through speech technology has predominantly utilized structured speech tasks (e.g., picture description), self-reported questionnaires, and controlled protocol-based interactions [4]. Recent studies have shown potential in analyzing voice interactions with smart speakers [5] and phone conversations [6], demonstrating associations with neuropsychological test scores. However, these approaches may not capture unstructured free-speech patterns that characterize everyday voice interactions. Furthermore, existing methods often analyze cross-sectional rather than longitudinal data, which is crucial for detecting the subtle progression of cognitive decline.
@@ II-A Detecting Cognitive Decline of Dementia
Cognitive decline detection research has primarily utilized the DementiaBank corpus for dementia detection that contains speech recordings from structured clinical interviews, such as picture description and fluency tasks [4]. These datasets present limitations: most contains cross-sectional data rather than longitudinal change tracking. Traditional approaches focus on identifying paralinguistic and linguistic markers indicative of dementia, including reduced word choices, increased pauses, reduced lexical diversity, and syntactic simplification. Recent natural language processing technologies enable automatic extraction of these features from speech transcripts, with transformer-based models such as BERT [7] demonstrating effectiveness in capturing semantic and lexical changes associated with dementia. However, these studies apply to the onset of dementia and Alzheimer’s disease rather than MCI, presenting challenges in early-stage detection.
@@ III-A Data Collection Details
We conducted an 18-month pilot study (2022-2024) with 35 older adults aged (16 females, 19 males; 61–80 years old, mean: 72, SD: 5.1) to collect VAS voice interactions for cognitive decline detection. It underwent seven session-based assessments and daily in-home voice interactions. A subset of 15 participants (7 females, 8 males; age mean: 74) consented to daily in-home voice command collection.
@@ III-A2 At-Home Voice Commands
The 15 participants interacted with Alexa daily, with recordings and transcripts stored in Amazon’s secure portal. All other data, such as consents, interviews, demographics, and MoCA scores, are stored in the HIPAA-compliant REDCap database. In the initial session, the RA provided setup training. The RA monitored participants’ device usage and gave email support when participation significantly decreased. On average, each participant generated approximately 47 commands weekly.
@@ IV-A Problem Formulation and Notation
We study cognitive status assessment using a longitudinal dataset of participants, each participant with a cognitive label determined by clinical MoCA evaluation. Our data comprises voice interactions collected over 18 months, indexed as . For each participant at month , we collect voice commands and their corresponding transcripts , from which we derive acoustic features and linguistic features . These are concatenated to form multimodal representations:
The complete time-series sequence of multimodal features for participant is denoted as , where each is month ’s multimodal features.
@@ V-A2 Iterative Prompt Optimization Details
Our algorithm begins with an initial prompt and performs automatic prompt optimization using a minibatch of training () and validation data () from monthly transcripts. The iterative refinement runs for three iterations, based on evidence that 3–5 iterations can yield effective prompts [12]. The optimized prompt improves the extraction of linguistic features relevant to MCI detection. Specifically, the process consists of:
@@
where denotes the F1-score evaluation metric. The optimized prompt is used to extract linguistic features and generate summaries for these features that are then fed into SBERT.
@@ V-B1 Speech Feature Embeddings
We employ the pre-trained HuBERT model [13] to extract acoustic feature embeddings from monthly command audio. HuBERT is trained via self-supervision on large-scale speech corpora to capture speaker characteristics, including prosody (pitch, rhythm), spectral patterns, and articulation relevant to MCI detection. Formally, let be the HuBERT feature extractor. For each audio command , we obtain an acoustic embedding: .
@@ V-B3 Temporal Sequence Modeling with Transformer
We use a time-series model to capture temporal patterns in the sequence of historical data from months 1 to (). The model takes as input the sequential multimodal embeddings , where each represents the multimodal features from month : .
We implement and compare transformer-based models such as PatchTST [14] and iTransformer [15], which have shown strong performance in modeling time-series data in several domains. The model takes as input and outputs a cognitive status prediction: , where denotes the learnable parameters of the transformer.
@@ VII-A Comparison with Baseline Models
Effectiveness of Cog-TiPRO: Cog-TiPRO consistently improves performance across most LLM variants. When combined with GPT4o-mini and iTransformer, it achieves the highest performance (73.80% acc, 72.67% F1), showing a 19.87% improvement over its baseline (53.93% acc). GPT3.5 with iTransformer ranks second (70.00% acc, 66.90% F1). These results indicate that advanced LLMs are more effective at identifying MCI-relevant aspects and patterns in command usage when guided by our iterative prompt refinement. This demonstrates the strength of our LLM-driven strategy. Additionally, multimodal fusion outperforms single-modality baselines, suggesting that feature fusion provides complementary information to enhance feature quality and detection accuracy.
@@ VIII Limitations
Our study has several limitations. First, despite 18 months of longitudinal data, our pilot study includes a small sample size. We plan to expand the number of participants in future work to apply Cog-TiPRO to broader and more diverse cohorts. Second, variability in home environments and reliance on voice commands may impact the quality of features related to cognitive decline and affect model performance. Our future studies will incorporate task designs and data collection that better capture established cognitive decline markers to improve prediction accuracy. Furthermore, we will address participants’ concerns regarding data privacy, usage, and bias in future studies. Lastly, our iterative prompt optimization increases computational overhead and inference time, which is a challenge for deployment in low-resource settings and real-time monitoring. We will study this problem in future work.
@@ IX Conclusion
In this paper, we demonstrate the viability of using VAS for the longitudinal collection and analysis of voice commands to detect cognitive decline. By collecting daily interactions from 15 older adults over 18 months in their homes, we capture speech patterns that reflect cognitive functioning through our Cog-TiPRO framework. Cog-TiPRO leverages LLMs to extract linguistic features indicative of MCI from unstructured, short, and noisy voice commands. The integration of LLM-driven prompt refinement, acoustic feature extraction, and temporal modeling enables more accurate detection of MCI. Using iTransformer, our approach achieves 73.80% accuracy and 72.67% F1-score, outperforming its baseline by 27.13%. The linguistic features extracted not only reveal patterns specific to everyday VAS interactions but also align with established clinical markers of MCI, demonstrating the potential of passive monitoring for early cognitive decline detection.

### v3

@@ Abstract
Early detection of cognitive decline is crucial for enabling interventions that can slow neurodegenerative disease progression. Traditional diagnostic approaches rely on labor-intensive clinical assessments, which are impractical for frequent monitoring. Our pilot study investigates voice assistant systems (VAS) as non-invasive tools for detecting cognitive decline through longitudinal analysis of speech patterns in short and unstructured voice commands. Over an 18-month period, we collected voice commands from 35 older adults, with 15 participants providing daily at-home VAS interactions. To address the challenges of analyzing these short, unstructured and noisy commands, we propose Cog-TiPRO, a framework that combines (1) LLM-driven iterative prompt refinement for linguistic feature extraction, (2) HuBERT-based acoustic feature extraction, and (3) transformer-based temporal modeling. Using iTransformer, our approach achieves 73.80% accuracy and 72.67% F1-score in detecting MCI, outperforming its baseline by 27.13%. Through our LLM approach, we identify linguistic features that uniquely characterize everyday command usage patterns in individuals experiencing cognitive decline.
@@
Mild Cognitive Impairment (MCI), a precursor to dementia, is characterized by subtle declines in cognitive functions such as attention, memory, and language. Early detection of these changes enables timely intervention to mitigate progression to dementia. However, detection and continuous monitoring of MCI remain challenging due to the absence of gold standard biomarkers [1]. Additionally, determining whether cognitive changes represent pathological decline requires complex clinical judgment. Currently, cognitive impairment often remains undetected until significant progression to dementia has occurred, leaving patients and families unprepared for managing its cognitive and functional impacts [2, 3].
Traditional diagnostics for MCI rely on clinical assessments, including neuroimaging, laboratory tests, and standardized neuropsychological examinations such as the Montreal Cognitive Assessment (MoCA) [4]. While effective for diagnosis, these methods have limitations: They require administration by trained healthcare professionals in clinical settings, making them labor-intensive and expensive for frequent monitoring. Since speech impairments such as increased pauses, word-finding difficulty, and language disfluency are among the earliest symptoms of cognitive decline, speech processing represents a promising approach for non-invasive and cost-effective MCI detection that could potentially be implemented in the everyday environment [5].
Prior research in cognitive monitoring through speech technology has predominantly utilized structured speech tasks (e.g. picture descriptions) [6, 7, 8], self-reported questionnaires, and controlled protocol-based interactions [9]. Recent studies have shown potential in analyzing voice interactions with smart speakers [10] and phone conversations [11], demonstrating associations with neuropsychological test scores. However, these approaches may not capture unstructured free-speech patterns that characterize everyday voice interactions. Furthermore, existing methods often analyze cross-sectional rather than longitudinal data, which is crucial for detecting the subtle progression of cognitive decline.
@@ II-A Detecting Cognitive Decline of Dementia
Cognitive decline detection research has primarily utilized the DementiaBank corpus for dementia detection that contains speech recordings from structured clinical interviews, such as picture description and fluency tasks [9]. These datasets present limitations: most datasets contain cross-sectional data rather than longitudinal change tracking. Traditional approaches focus on identifying paralinguistic and linguistic markers indicative of dementia, including reduced word choices, increased pauses, reduced lexical diversity, and syntactic simplification. Recent natural language processing (NLP) technologies enable automatic extraction of these features from speech transcripts, with transformer-based models such as BERT [12] demonstrating effectiveness in capturing semantic and lexical changes associated with dementia. However, these studies apply to the onset of dementia and Alzheimer’s disease rather than MCI, presenting challenges in early-stage detection.
@@ III-A Data Collection Details
We conducted an 18-month pilot study (2022-2024) with 35 older adults aged (16 females, 19 males; 61–80 years old, mean: 72, SD: 5.1) to collect VAS voice interactions for cognitive decline detection. Participants were recruited through flyers, with informed consent obtained and procedures approved by the Institutional Review Board (IRB) to ensure ethical standards. Data collection underwent seven session-based assessments and daily in-home voice interactions. A subset of 15 participants (7 females, 8 males; age mean: 74) consented to daily in-home voice command collection. The remaining 20 participants contributed only session-based data. We utilized the 15 participants’ in-home longitudinal data in further analysis.
@@ III-A2 At-Home Voice Commands
The 15 participants interacted with Alexa daily at home, with recordings and transcripts stored in Amazon’s secure portal. All other data, such as consents, interviews, demographics, and MoCA scores, are stored in the HIPAA-compliant REDCap database. In the initial session, the RA provided setup training. The RA monitored participants’ device usage and gave email support when participation significantly decreased. On average, each participant generated approximately 47 commands weekly.
@@ IV-A Problem Formulation and Notation
We study cognitive status assessment using an in-home longitudinal dataset containing participants, each participant with a cognitive label determined by the clinical MoCA evaluation. Our data comprises voice interactions collected over months, indexed as . For each participant at month , we collect voice commands and their corresponding transcripts , from which we derive acoustic features and linguistic features . These are concatenated to form the multimodal representations:
The complete time-series sequence of multimodal features for participant is denoted as:
where each is month ’s multimodal features.
@@ V-A2 Iterative Prompt Optimization Details
Our algorithm begins with an initial prompt and performs automatic prompt optimization using a minibatch of training () and validation data () from monthly transcripts. Using a small and diverse minibatch gives fast and low-cost feedback in short cycles that can guide effective edits without overfitting [17]. The iterative refinement runs for three iterations, based on evidence that 3–5 iterations can yield effective prompts [18]. The optimized prompt improves the extraction of linguistic features relevant to MCI detection. Specifically, the process consists of:
@@
where denotes the F1-score evaluation metric. Optimized prompt is used to extract linguistic features and generate summaries for these features that are then fed into SBERT.
@@ V-B1 Speech Feature Embeddings
We employ the pre-trained HuBERT model [19] to extract acoustic feature embeddings from monthly command audio. HuBERT is trained via self-supervision on large-scale speech corpora to capture speaker characteristics, including prosody (e.g. pitch, rhythm, pause), spectral patterns, and articulation relevant to MCI detection. Formally, let be the HuBERT feature extractor. For each audio command , we obtain its acoustic embedding:
@@ V-B3 Temporal Sequence Modeling with Transformer
We use a time-series model to capture temporal patterns in the sequence of historical data from months 1 to (). The model takes as input the sequential multimodal embeddings:
where each is the multimodal features from the month : .
We implement and compare transformer-based models such as PatchTST [20] and iTransformer [21], which have shown strong performance in modeling time-series data in several domains. The model takes as input and outputs a cognitive status prediction:
where denotes the learnable parameters of the transformer.
The complete Cog-TiPRO can be found in Algorithm 1.
@@ VII-A Comparison with Baseline Models
Effectiveness of Cog-TiPRO: Cog-TiPRO consistently improves performance across most LLM variants. When combined with GPT4o-mini and iTransformer, it achieves the highest performance (73.80% acc, 72.67% F1), showing a 19.87% improvement over its baseline (53.93% acc). GPT3.5 with iTransformer ranks second (70.00% acc, 66.90% F1). These results indicate that advanced LLMs are more effective at identifying MCI-relevant aspects and patterns in command usage when guided by our iterative prompt refinement. This demonstrates the strength of our LLM-driven strategy with pre-trained LLM models without the need for specific fine-tuning, in contrast to previous work [27]. Additionally, multimodal fusion outperforms single-modality baselines, suggesting that feature fusion provides complementary information to enhance feature quality and detection accuracy.
@@ VIII Limitations
Our study has several limitations. First, despite 18 months of longitudinal data, our pilot study includes a small sample size. We plan to expand the number of participants in the future work to apply Cog-TiPRO to broader and more diverse cohorts. Second, variability in home environments and reliance on voice commands may impact the quality of features related to cognitive decline and affect model performance. Our future studies will incorporate task designs and data collection that better capture established cognitive decline biomarkers to improve prediction accuracy and reliability. Furthermore, we will address participants’ concerns regarding data privacy, usage, and bias in future studies. Lastly, our iterative prompt optimization increases computational overhead and inference time, which is a challenge for deployment in low-resource settings and real-time monitoring. We will focus on this problem in our future study.
@@ IX Conclusion
In this paper, we demonstrate the viability of using VAS for the longitudinal data collection and analysis of voice commands to detect cognitive decline. By collecting daily interactions from 15 older adults over 18 months in their homes, we capture speech patterns that reflect cognitive functioning through our Cog-TiPRO framework. Cog-TiPRO leverages LLMs to extract linguistic features and aspects indicative of MCI from unstructured, short, and noisy voice commands. The integration of LLM-driven iterative prompt refinement, acoustic feature extraction, and temporal modeling enables more accurate detection of MCI. Using iTransformer, our approach achieves 73.80% accuracy and 72.67% F1-score, outperforming its baseline by 27.13%. The linguistic features extracted not only reveal patterns specific to everyday VAS interactions but also align with established clinical markers of MCI, demonstrating the potential of passive monitoring for the early detection of cognitive decline.
X Acknowledgement
This research is funded by the US National Institutes of Health National Institute on Aging R01AG067416. We thank Professor David Kotz and Professor Brian MacWhinney for providing valuable guidance.